This is the Title of the Thesis Broken Across Two Lines

by

Author’s Name Here

(Under the direction of Professor’s Name Here)

Abstract

This is the abstract, a brief summary of the contents of the thesis. It is limited to 150 words in length.

The abstract page(s) are not numbered and are not necessarily included in a printed and bound copy of the thesis. The index words at the bottom of the abstract should be chosen carefully, preferably with the help of your colleagues.

Index words: Index word or phrase, Index word or phrase, Index word or phrase, Index word, Index word, Index word

by

Author’s Name Here

B.A., University of Somewhere, 1998 M.A., University of Somewhere Else, 2001

A Dissertation Submitted to the Graduate Faculty of The University of Georgia in Partial Fulfillment

of the Requirements for the Degree

Doctor of Philosophy

Athens, Georgia

\begin{matrix}c \\ ◯ \end{matrix}2010

Author’s Name Here This is the Title of the Thesis

Broken Across Two Lines

by

Author’s Name Here

Approved:

Major Professors: Professor’s Name Professor’s Name

Committee: Member’s Name Member’s Name Member’s Name

Electronic Version Approved:

Dean’s Name Here Dean of the Graduate School The University of Georgia June 2010

Author’s Name Here June 29, 2010

Acknowledgments

You can have acknowledgments or a preface here.

Contents

1 This is the First Chapter 5

1.1 This is a section in the first chapter . . . . . . . . . . . . . . . . . . . . . . . 5 2 What L^{A}TEX is all about 7

List of Figures

1.1 Example of a figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

List of Tables

1.1 Example of a table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Chapter 1

This is the First Chapter

1.1 This is a section in the first chapter

This is an example of how to format a thesis with LaTeX in the simplest possible way. This is a simple example. This is a simple example and will be very boring to read. This is an example of how to format a thesis with LaTeX in the simplest possible way.

This is a simple example. This is a simple example and will be very boring to read. This is an example of how to format a thesis with LaTeX in the simplest possible way. This is a simple example. This is a simple example and will be very boring to read. This is an example of how to format a thesis with LaTeX in the simplest possible way.

This is a simple example. This is a simple example and will be very boring to read. This is an example of how to format a thesis with LaTeX in the simplest possible way. This is a simple example. This is a simple example and will be very boring to read. This is an

[ You could put a picture here. ] Figure 1.1: Example of a figure.

Table 1.1: Example of a table. [The contents of the table would go here.]

example of how to format a thesis with LaTeX in the simplest possible way.

This is a simple example. This is a simple example and will be very boring to read. This is an example of how to format a thesis with LaTeX in the simplest possible way. This is a simple example. This is a simple example and will be very boring to read.

Proceed with more sections and chapters, bibliography, etc., after this.

Chapter 2

What L^{A}TEX is all about

This document is an example of how to format a thesis or dissertation using L^{A}TEX and get results acceptable at The University of Georgia.

L^{A}TEX (with its parent TEX) has two major advantages for academic use. First, to a remarkable degree it makes design decisions automatically. The author supplies only the words of a text, and L^{A}TEX places them on the page in an aesthetic manner, avoiding rivers and awkward breaks. In this respect LaTeX is like a very intelligent typist or typesetter.

Second, L^{A}TEX can typeset complex mathematical formulas such as

∑_{i=1}^{∞} x^{y+z} = \frac{p + q + r}{s + t + u + v}

both displayed (as shown above) and in the text, as in ∑_{i=1}^{∞} x^{y+z} = \frac{p+q+r}{s+t+u+v}. This makes TEX and L^{A}TEX indispensable for mathematicians, physicists, and the like.

LaTeX also has built–in formats for other kinds of displayed material such as verse,

Freude, schöne Götterfunken Tochter aus Elysium, Wir betreten, feuertrunken, Himmlische, dein Heiligtum! Deine Zauber binden wieder Was der Mode streng geteilt…

and direct quotations:

The society that scorns excellence in plumbing, because plumbing is a humble activity, and tolerates shoddiness in philosophy, because philosophy is an exalted activity, will have neither good plumbing nor good philosophy. Neither its pipes nor its theories will hold water.

— John Gardner, Excellence

If you wish, quotes and other displayed material can be single–spaced; here is an example of how that is achieved:

Yes, I wrote “The Purple Cow.” I’m sorry now I wrote it. But I can tell you anyhow I’ll kill you if you quote it!

— Anonymous?

Well, maybe it’s not as anonymous as it looks.

Computer scientists use LaTeX’s “verbatim” format to display portions of computer programs in text, like this:

1   GO TO 2
2   GO TO 1
3   PRINT "THIS STATEMENT WILL NEVER EXECUTE"
4   END

There you have it.^{1}

1This is a footnote. Notice that footnotes are single–spaced even though the text is double–spaced. Long footnotes are discouraged; either make important points in the text or leave them out.

Whenever you quote parts of a computer program in English text, they should be set off by using typewriter type. The PRINT statement in the program above will never execute because the GO TO statement above it keeps execution from reaching it.